In the cross aaBbcc X AaBbCc, what would be the probability that the offspring has a phenotype identical to either one of the parents?
Answer #1

Answer:

The probability that the offspring has a phenotype identical to either one of the parents = 1/8

Explanation:

aaBbcc X AaBbCc --Parents

aa x Aa = Aa (1/2) & aa (1/2)

Bb x Bb = BB (1/4), Bb (1/2) & bb(1/4)

cc x Cc = Cc (1/2) & cc (1/2)

Probability of aaBbcc = 1/2 * 1/2 * 1/2 = 1/8

Probability of AaBbCc = 1/2 * 1/2 * 1/2 = 1/8